OpenStudy (anonymous):

medals Tom and Jerry enter a race. Tom runs an average of 10 kilometers per hour, and Jerry runs an average of 9 kilometers per hour. If Tom finishes one half hour before Jerry, how far do Tom and Jerry run? 20 km 90 km 45 km 22.5 km

OpenStudy (perl):

since they are both finishing a race, lets call the distance of the race d. d = 10 (t-1/2) d= 9*t 10(t-1/2)= 9t 10t - 5 = 9t t = 5 . so d= 9*5 = 45 they both ran 45 kilometers

OpenStudy (perl):

unless i missed something, medals please :/

OpenStudy (jack1):

puggles are awesome.. def medals

OpenStudy (perl):

hey thats a platypus, not a puggle

OpenStudy (jack1):

o.0 not knowing what's in ur own pic... def worth subtracting medals

OpenStudy (jack1):

-.- puggle: juvenile platypus or echidna